OpenStudy (anonymous):

Jerry invested $14,000 in a certificate of deposit at 5%. He also put $2,500 in a savings account at 3%. How much interest will he earn after one year?

OpenStudy (nurali):

$775 interest earned.

OpenStudy (anonymous):

ok thank u bunches

OpenStudy (mathstudent55):

5% of 14,000 = 0.05 * 14,000 = 700 3% of 2500 = 0.03 * 2500 = 75 700 + 75 = 775